# FactTrack: Time-Aware World State Tracking in Story Outlines

### v1

@@ 6 Ablation Study
    Replace NLI Model in Detect Contradictions with Other Models: Our motivation for using the NLI model is to reduce the computational cost of the entire pipeline. We explore replacing the NLI model with few-shot LLMs and provide a quantitative estimation of the improvement.
6.1 Results
@@ 7 Conclusion and Future Work
Appendix A Methodology Details
A.1 Prompt of event Decomposition
@@
A.2 Epsilon Padding on Timeline
Note that to easily distinguish the boundary between different events, we set a tiny number $\epsilon=10^{-6}$ as the padding between the events. So the real format in of event time interval is:
@@ A.2 Epsilon Padding on Timeline
A.3 Pseudocode of Interval Operation
In our data structure 444 Refer to this file for the core part of our data structure: https://github.com/cogito233/contradict-detector/blob/main/contradict-detector/core/contradict_detector.py. , we use a list to store the intervals of all fact content, their validity intervals, and their embeddings from the contriever model. Here are some definitions of the notation in the pseudocode:
@@ A.3 Pseudocode of Interval Operation
    For $Same(p,P)$ as considered in the update condition, we drop identical atomic facts to reduce computation, as judged by $sim_{contriver}(p,P)>0.95$.
@@
    For $Contradict(p,P)$, we use different thresholds to make detection more effective and make the update more robust. In the determine and update validity interval step, this means $NLI\_Score>0.8$, while in check contradict step, this means $NLI\_Score>0.2359$, set as discussed in Appendix A.4.
@@
A.4 Finetuning NLI Models
Since the task format of recognizing whether two atomic facts contradict is similar to NLI, we use an NLI model from huggingface (MoritzLaurer/DeBERTa-v3-large-mnli -fever-anli-ling-wanli) and finetune it using GPT-4-Turbo annotations to adapt the narration domain, see the prompt in Table 5. To construct the dataset, we generate 60 outlines from the WritingPrompts dataset Fan et al. (2018) and use 50 for training and 10 for testing. In the training set, we have a total of 856,748 fact pairs. To create a subsample, we first use an NLI model without any finetuning to retrieve data. Then, we employ a second NLI model, which has been finetuned based on the data retrieved by the first, non-finetuned model. From the output of both models, we subsample 10,000 fact pairs based on each model. As the vast majority of fact pairs are not contradictory, to improve class balance, for each model, we randomly select 5k from the top 1% most confidently predicted contradictions; 2k from the top 1-10%; and 3k from the remaining. After removing duplicate data, we get 18,702 fact pairs, which are then annotated by GPT-4-Turbo using the prompt shown in Table 5.
Based on this annotation result, we estimate the overall positive rate of all 856,748 fact pairs as $2.98\%$. To evaluate, we re-rank the fact pairs in the test set, and divide the top 10% of data into 10 intervals, with 100 data points randomly sampled in each interval to observe the positive rate. Therefore, we set the percentile threshold for flagging a contradiction in FactTrack at 3% ($NLI\_Score\geq 0.2359$), with precision = 60% and recall = 60% on our GPT-4-Turbo-annotated test set.
@@ A.4 Finetuning NLI Models
Table 5: The prompt given to GPT-4-Turbo to classify two fact statements as contradicting each other. We use the result to fine-tune our NLI model to better fit the narrative domain.
A.5 Outline Generation
In our experiments, we use depth=3, branching factor=3 to generate outlines, where depth means the maximum number of hierarchical outline layers, and the branching factor means the number of sub-events one node decomposes into. We chose these settings because the length of such an outline (around 4800 tokens for GPT-4-Turbo) makes the outline sufficiently complex while still somewhat tractable for annotation. FactTrack itself can easily scale to a more complex outline. Compared with the Detailed Outliner in the DOC paper, we have two changes. (1) We add the sections “begin event” and “end event” to each event, because we want the Outline to be richer in facts and more precise in context; (2) We set the “partial outline” (prompt context for generating each new sub-event) to always be the content’s direct parent, while in Detailed Outliners, the default setting is to be all ancestors and their siblings. This is mainly because we consider our method as using constant text window size for outline generation. Additionally, in our code (to be open-sourced upon publication), we do implement all the options to conveniently change the above settings, which we believe will facilitate further evaluation for new methods in the future. Our prompts are shown in Table 6.
@@ A.5 Outline Generation
Table 6: The prompt given to GPT-4-Turbo to generate the outline. Compared to DOC, we added the boundary events to make the generation more stable and have a higher fact density.
Appendix B Baselines and Evaluation Settings
B.1 Human Annotation Attempt
We initially attempted human annotation but found the task to be highly challenging to annotate. Even experienced human labor and advanced LLMs struggled to accurately obtain ground truth labels. This situation has become increasingly common in the era of LLMs. We used prolific.co as our human annotation platform and implemented a foldable outline with the Potato Annotation Framework Pei et al. (2022), along with some structured input formats (See Figures 6 and 7). The budget for each annotator was 30 minutes per Outline, at 15 USD per hour. The prompts given to human annotators were also used to establish a Full Context Detection Baseline for GPT-4, which can be referred to in Table 8. In this context, we encountered problems such as: (1) annotators struggling to understand the form of the question; (2) 30 minutes being insufficient for annotators to retrieve all ground truth labels; (3) some annotators attempting to cheat by using GPT-4 for annotation. It is worth noting that it is theoretically feasible for human annotators to complete all the annotation tasks mentioned in this text. However, after our initial exploration, we concluded that sufficiently reducing annotation noise and achieving statistically significant conclusions would result in astonishing economic costs and probability infeasible. Therefore, in this project, we use the most advanced version of GPT-4-Turbo to complete all annotation tasks.
We in fact found that using GPT-4-Turbo, with the same instructions as a human annotator, achieved better results compared to humans upon manual inspection. However, we also found on inspection that even when our method FactTrack was deployed on LLaMA2-7B-Chat, it maintained similar or even higher quality than GPT-4-Turbo. Using a worse baseline to evaluate our method was unreasonable, so we ultimately decided to use a scoring-based metric that considers individual pairs of facts rather than looking at full outlines. (We also tried using GPT-4-Turbo for preference annotation but found it to be noisy. This was due to different events potentially contradicting in various aspects, making it difficult to determine which was “better” or “worse.”)
@@ B.1 Human Annotation Attempt
B.2 Prompts Used in Baseline
@@
B.3 Prompts Used in Metrics
@@
B.4 Prompts Used in Ablation
@@
Appendix C Experiment Statistics
We used 100 prompts to generate 90 Outlines from the writing prompt, with statistics in Table 12. For FactTrack deploy on LLaMA2-7B-Chat, we get 4589 positive pairs in total; For FactTrack deploy on GPT-4-Turbo, we get 8883 positive pairs in total; Details shown in Table 13. For Full Outline Detection on GPT-4-Turbo, we get 408 positive pairs in total; for Pairwise Detection on LLaMA2-7B-Chat, we subsample 10% pairs and get 400 positive pairs, means if we run the whole population, we expect get around 4,000 positive pairs of this baseline. Since we can not easily change the thresholding for our baseline, to make the comparison fairer, we subsample our method’s predictions so that both methods flag a similar total number of contradictions. To compare with Full Outline Detection on GPT-4-Turbo with 298 positive pairs detected, we subsample by the Max fact pair NLI Score between event pairs for 300 samples, while comparing with Full Outline Detection on GPT-3.5-Turbo (829 positive pair), Pairwise Detection on LLaMA2-7B-Chat (4000 positive pairs in estimation), we execute pure random sampling for 500 samples as a representative of the whole population (4589 positive pairs) of our methods.
Tables 14 and 15 show the sample size and distribution of our evaluation set. Several data points fail the generation, and we simply drop them. By observing the distribution, we find that FactTrack’s detection capability is slightly inferior to GPT-4-Turbo on clear contradictions (5 score), which may also be because we subsample top 300 instead of top 298, but on more nuanced or partial contradictions (2 score, 3 score, 4 score) FactTrack is significantly better than the two baselines. This observation intuitively confirms the source of strength for our approach.
@@ Appendix C Experiment Statistics
Table 12: Statistics of outlines.
@@
# Total Contradictions: 4589 8883
Positive Rate: 0.07 0.14
@@
Layer 1: 350 667
Layer 2: 1140 2170
Layer 3: 2943 5801
Layer 1 $\leftrightarrow$ Layer 2: 159 171
Layer 2 $\leftrightarrow$ Layer 1: 404 774
Layer 1 $\leftrightarrow$ Layer 3: 457 624
Layer 2 $\leftrightarrow$ Layer 3: 1074 2193
Table 13: Contradictions detected for FactTrack LLaMA2-7B and FactTrack GPT-4.
@@
Full Outline Detection (GPT-4-Turbo) 298 296 35.14% 34.12% 15.20% 9.46% 6.08%
FactTrack (GPT-4-Turbo, top 300) 300 300 4.33% 29.67% 25.00% 30.33% 10.67%
@@
Table 14: Distribution of Pairwise Score for different methods. Magnitude means across 90 outlines, how many postive samples are detected by a given method. Sample size is the number of randomly selected samples we let GPT-4-Turbo to annotate.
@@
Full Outline Detection (GPT-4-Turbo) 298 298 34.90% 23.49% 14.77% 13.42% 13.42%
FactTrack (GPT-4-Turbo, top 300) 300 299 21.07% 37.13% 14.72% 15.05% 12.04%
@@
Table 15: Distributon of Context Score. Magnitude means out of 90 outlines, how many postive samples were detected by a given method. Sample size is the number of randomly selected samples we let GPT-4-Turbo annotate.
Appendix D Using FactTrack to Enhance Outline Generation
We also implement algorithms to enhance outline generation itself by both direct rewriting, rewriting conditional on facts (see Table 16), and rewriting conditional on events (see Table 17). The qualitative inspection shows that it can improve factual consistency since we can detect and solve the problem on high-level planning. Sometimes our method can make false positive predictions, and continuing rewrite, we apply one of the following two strategies: (1) rewrite the event until rewrite sampling time (by default 5 times) and then ignore it; (2) resample the event until maximum sampling time (by default 10 times) and then rewrite the whole outline. Both strategies can help generate a high-quality outline.

### v2

@@ 5.1 Ablation Study
    Replacing the NLI Model in Detecting Contradictions: We use the NLI model to lower the pipeline’s computational cost. We explore replacing it with few-shot LLMs and estimate the potential improvements quantitatively.
5.1.0.0.1 Results
@@
5.2 Document-level contradict detection
We extended the application of FactTrack to the ContraDoc dataset Li et al. (2023), a dataset for evaluating document-level contradictions. In alignment with the original study, we employed a binary judgment framework, which involves directly prompting large language models (LLMs) as a baseline. Our approach, FactTrack, demonstrated a better F1-score compared to the baseline models. The observed reduction in precision relative to the baseline can be attributed to the presence of borderline contradictions within the documents and the ambiguity in the decomposition process. The results are detailed in Table 3.
Experiment Precision Recall F1
GPT4o-mini 89.29% 5.57% 10.48%
FactTrack 52.71% 62.81% 57.32%
Table 3: Performance of the baseline and FactTrack in the Binary Judgement experiment in ContraDoc. FactTrack was implemented using GPT4o-mini, with splitting events by sentences conducted via the nltk package.
@@ 6 Conclusion and Future Work
Limitations
The difficulty of identifying all contradictions and partial contradictions significantly constrained our evaluations (see Appendix C.1), limiting us to obtaining gold labels of contradictions. Therefore, we used GPT-4 to annotate the contradicting score as a proxy for the evaluation task.
The context window size of baseline and evaluation metrics also restricted us from running experiments on outlines much longer than the current 2000-3000 words. While FactTrack is capable of detecting contradictions in such outlines with near-linear cost growth, we did not evaluate them in this work due to the potential performance degradation of GPT-4 in longer contexts.
Additionally, the challenge of conducting thorough evaluations impacted the system’s development. Many decisions, such as prompt design, the choice of models for each module (e.g., the few-shot language model outperforms the NLI model as shown in Table 2), and the selection of hyperparameters (e.g., thresholds for the contriver and NLI models), were made manually rather than through rigorous validation. As a result, there may be considerable room for improvement in the detailed design of individual modules.
Finally, FactTrack’s performance may decrease on LLMs that lack strong generation and instruction-following capabilities.
@@
Appendix A Concept Clarification
A.1 Time-aware Contradiction
A time-aware contradiction occurs when two factual statements remain contradictory even when temporal context is considered. For instance, the statements:
  • “Donald Trump is the president of the United States.”
  • “Joe Biden is the president of the United States.”
may appear contradictory. However, when interpreted with temporal context—Trump serving before 2021 and Biden serving from 2021 onward—the contradiction is resolved. Thus, this is not a time-aware contradiction. Instead, a time-aware contradiction persists even after considering temporal information.
A.2 Time-dependent Fact
A time-dependent fact is a fact that holds true within a specific time interval but may become invalid afterward. For example:
“Joe Biden is the president of the United States from 2021 to 2024.”
is a time-dependent fact since its validity is constrained to a specific period. In our FactTrack setting, we maintain a relative timeline based on event order rather than absolute timestamps, allowing for contextual fact verification without requiring explicit temporal annotations.
A.3 Dynamic Nature of the World State
The dynamic nature of the world state refers to the updates in the state of entities following an event. Consider the following scenario:
“After buying the book, Eva leaves the store.”
This event results in state changes:
  • Eva’s location: “Eva is in the store” $\rightarrow$ “Eva is not in the store.”
  • Book ownership: “Eva does not own the book” $\rightarrow$ “Eva owns the book.”
Capturing such transitions is essential for factuality assessment, as ignoring these updates can lead to incorrect conclusions about entity states.
Appendix B Methodology Details
B.1 Prompt of event Decomposition
See Table 4 for more detail.
@@ B.1 Prompt of event Decomposition
B.2 Epsilon Padding on Timeline
Note that to easily distinguish the boundary between different events, we set a tiny number $\epsilon=10^{-6}$ as the padding between the events. Therefore, the real format of the event time interval is:
@@ B.2 Epsilon Padding on Timeline
B.3 Pseudocode of Interval Operation
In our data structure 444 Refer to this file for the core part of our data structure: https://github.com/cogito233/fact-track/blob/main/fact-track/core/contradict_detector.py. , We use a list to store the intervals of all fact content, their validity intervals, and their embeddings, using Contriever (Izacard et al., 2021) as a retrieval model to filter out irrelevant fact pairs. Here are some definitions of the notation in the pseudocode:
@@ B.3 Pseudocode of Interval Operation
    For $Same(p,P)$, as considered in the update condition, we drop identical atomic facts to reduce computation, as judged by $sim_{contriver}(p,P)>0.95$.
@@
    For $Contradict(p,P)$, we use different thresholds to make detection more effective and to make the update more robust. In the determine and update validity interval step, this means $NLI_{S}core>0.8$, while in the check contradiction step, this means $NLI_{S}core>0.2359$, as discussed in Appendix B.4.
@@
B.4 Finetuning NLI Models
Since the task of recognizing whether two atomic facts contradict is similar to NLI, we use an NLI model from Hugging Face (MoritzLaurer/DeBERTa-v3-large-mn li-fever-anli-ling-wanli) and fine-tune it using GPT-4 annotations to adapt to the narrative domain. See the prompt in Table 5. To construct the dataset, we generate 60 outlines from the WritingPrompts dataset Fan et al. (2018) and use 50 for training and 10 for testing. In the training set, we have a total of 856,748 fact pairs. To create a subsample, we first use an NLI model without any finetuning to retrieve data. Then, we employ a second NLI model, which has been finetuned based on the data retrieved by the first, non-finetuned model. From the output of both models, we subsample 10,000 fact pairs for each model. As the vast majority of fact pairs are not contradictory, to improve class balance, we randomly select 5k from the top 1% most confidently predicted contradictions for each model; 2k from the top 1-10%; and 3k from the remaining. After removing duplicate data, we get 18,702 fact pairs, which are then annotated by GPT-4 using the prompt shown in Table 5.
Based on this annotation result, we estimate the overall positive rate of all 856,748 fact pairs as $2.98\%$. To evaluate, we re-rank the fact pairs in the test set, and divide the top 10% of data into 10 intervals, with 100 data points randomly sampled in each interval to observe the positive rate. Therefore, we set the percentile threshold for flagging a contradiction in FactTrack at 3% ($NLI\_Score\geq 0.2359$), with precision = 60% and recall = 60% on our GPT-4-annotated test set.
@@ B.4 Finetuning NLI Models
Table 5: The prompt given to GPT-4 to classify two fact statements as contradicting each other. We use the result to fine-tune our NLI model to better fit the narrative domain.
B.5 Outline Generation
In our experiments, we use depth=3 and branching factor=3 to generate outlines, where depth refers to the maximum number of hierarchical outline layers, and branching factor refers to the number of sub-events into which one node decomposes. We chose these settings because the length of such an outline (around 4800 tokens for GPT-4) makes it sufficiently complex while still somewhat tractable for annotation. FactTrack can easily scale to a more complex outline. Compared with the Detailed Outliner in the DOC paper, we have two changes. (1) We add the sections “begin event” and “end event” to each event, because we want the Outline to be richer in facts and more precise in context; (2) We set the "partial outline" (prompt context for generating each new sub-event) to always be the content’s direct parent, whereas in Detailed Outliners, the default setting is to include all ancestors and their siblings. This is mainly because we consider our method as using constant text window size for outline generation. Additionally, in our code (to be open-sourced upon publication), we have implemented all the options to conveniently change the above settings, which we believe will facilitate further evaluation of new methods in the future. Our prompts are shown in Table 6.
@@ B.5 Outline Generation
Table 6: The prompt was given to GPT-4 to generate the outline. Compared to DOC, we added boundary events to make the generation more stable and increase fact density.
Appendix C Baselines and Evaluation Settings
C.1 Human Annotation Attempt
We initially attempted human annotation but found the task to be highly challenging. Even experienced human annotators and advanced LLMs struggled to accurately obtain ground truth labels. This situation has become increasingly common in the era of LLMs. We used prolific.co as our human annotation platform and implemented a foldable outline with the Potato Annotation Framework Pei et al. (2022), along with some structured input formats (see Figures 6 and 7). The budget for each annotator was 30 minutes per Outline, at 15 USD per hour. The prompts given to human annotators were also used to establish a Full Context Detection Baseline for GPT-4, which is referred to in Table 8. In this context, we encountered problems such as: (1) annotators struggling to understand the format of the question; (2) 30 minutes being insufficient for annotators to retrieve all ground truth labels; and (3) some annotators attempting to cheat by using GPT-4 for annotation. It is worth noting that it is theoretically feasible for human annotators to complete all the annotation tasks mentioned in this text. However, after our initial exploration, we concluded that sufficiently reducing annotation noise and achieving statistically significant conclusions would result in astonishing economic costs and probability infeasible. Therefore, in this project, we use the most advanced version of GPT-4 to complete all annotation tasks.
We found that using GPT-4, with the same instructions as a human annotator, achieved better results compared to humans upon manual inspection. However, we also found on inspection that even when our method FactTrack was deployed on LLaMA2-7B-Chat, it maintained similar or even higher quality than GPT-4. Using a worse baseline to evaluate our method was unreasonable, so we ultimately decided to use a scoring-based metric that considers individual pairs of facts rather than full outlines. (We also tried using GPT-4 for preference annotation but found it to be noisy. This was due to different events potentially contradicting each other in various aspects, making it difficult to determine which was “better” or “worse.”)
@@ C.1 Human Annotation Attempt
C.2 Prompts Used in Baseline
@@
C.3 Prompts Used in Metrics
@@
C.4 Prompts Used in Ablation
@@
Appendix D Experiment Statistics
We used 100 prompts to generate 90 Outlines from the writing prompt, with statistics in Table 12. For FactTrack deployed on LLaMA2-7B-Chat, we get 4589 positive pairs in total; for FactTrack deployed on GPT-4, we get 8883 positive pairs in total; details are shown in Table 13. For Full Outline Detection on GPT-4, we get 408 positive pairs in total; for Pairwise Detection on LLaMA2-7B-Chat, we subsample 10% of the pairs and get 400 positive pairs, meaning if we run the whole population, we expect to get around 4,000 positive pairs for this baseline. Since we cannot easily change the threshold for our baseline, to make the comparison fairer, we subsample our method’s predictions so that both methods flag a similar total number of contradictions. To compare with Full Outline Detection on GPT-4, which detected 298 positive pairs, we subsample by the Max fact pair NLI score between event pairs for 300 samples. While comparing with Full Outline Detection on GPT-3.5-Turbo (829 positive pairs) and Pairwise Detection on LLaMA2-7B-Chat (estimated 4000 positive pairs), we perform pure random sampling for 500 samples as a representative of the whole population (4589 positive pairs) of our method.
Tables 14 and 15 show the sample size and distribution of our evaluation set. Several data points failed during generation, and we simply dropped them. By observing the distribution, we find that FactTrack’s detection capability is slightly inferior to GPT-4 on clear contradictions (score 5), which may also be because we subsampled the top 300 instead of the top 298. However, on more nuanced or partial contradictions (scores 2, 3, and 4), FactTrack is significantly better than the two baselines. This observation intuitively confirms the strength of our approach.
Given the label-wise result, it is also possible to estimate a precision and recall value of the binary metrics. According to Tables 14, if we consider scores 4 and 5 as true positives and scores 1 to 3 as false positives, and assume there are 450 gold contradictions in all 90 outlines, our baseline of GPT-4 has $TP=46$, $FP=252$, $FN=404$, with $P=0.154$ and $R=0.102$. FactTrack on GPT-4 has $TP=123$, $FP=177$, $FN=327$, with $P=0.410$ and $R=0.273$, using the results in Table 13. However, this is only a rough estimate because the annotation of GPT-4 is not necessarily equivalent to the gold label, and the estimation of the total contradictions also requires precise annotation.
@@ Appendix D Experiment Statistics
Table 12: Statistics of outlines. Each outline contains three levels (3+9+27=39). Word and sentence counts were determined using NLTK for tokenization. The variance in word count across different outlines has been calculated. Unique words were identified by simple deduplication after tokenization. On average, each unique word appears 6.5 times within an outline and 25.6 times across all 90 outlines.
@@
# Total Contradictions: 4559 8883
Positive Rate: 7% 14%
@@
Layer 1: 52 19% 44 16%
Layer 2: 313 10% 623 20%
Layer 3: 2229 7% 4462 15%
Layer 1 $\leftrightarrow$ Layer 2: 184 8% 303 13%
Layer 1 $\leftrightarrow$ Layer 3: 333 5% 702 10%
Layer 2 $\leftrightarrow$ Layer 3: 1448 7% 2749 13%
Table 13: Contradictions detected for FactTrack LLaMA2-7B and FactTrack GPT-4 by FactTrack with the settings in Appendix B.1. At the same NLI threshold, GPT-4 performs better at contradiction detection than LLaMA2-7B, capturing more potential contradictions due to fine-grained and accurate decomposition. We also visualize the frequency of contradictions occurring between different layers of the outline. Additionally, this provides insight into the higher-density contradictions that may occur in a higher-level outline.
@@
Full Outline Detection (GPT-4) 298 296 35.14% 34.12% 15.20% 9.46% 6.08%
FactTrack (GPT-4, top 300) 300 300 4.33% 29.67% 25.00% 30.33% 10.67%
@@
Table 14: Distribution of Pairwise Score for different methods. Magnitude refers to how many positive samples are detected by a given method across 90 outlines. Sample size is the number of randomly selected samples we allowed GPT-4 to annotate.
@@
Full Outline Detection (GPT-4) 298 298 34.90% 23.49% 14.77% 13.42% 13.42%
FactTrack (GPT-4, top 300) 300 299 21.07% 37.13% 14.72% 15.05% 12.04%
@@
Table 15: Distribution of Context Score. Magnitude refers to how many positive samples were detected by a given method out of 90 outlines. Sample size is the number of randomly selected samples we allowed GPT-4 to annotate.
Appendix E Using FactTrack to Enhance Outline Generation
We also implement algorithms to enhance outline generation by direct rewriting, rewriting conditional on facts (see Table 16), and rewriting conditional on events (see Table 17). Qualitative inspection shows that this approach can improve factual consistency since it allows us to detect and solve problems at a high level of planning. Sometimes our method can make false positive predictions, and continuing rewrite, we apply one of the following two strategies: (1) rewrite the event until rewrite sampling time (by default 5 times) and then ignore it; (2) resample the event until maximum sampling time (by default 10 times) and then rewrite the whole outline. Both strategies can help generate a high-quality outline.